# Optical aberrations in autonomous driving: Physics-informed parameterized temperature scaling for neural network uncertainty calibration

### v1

@@ 3 Theoretical essentials
In this section, we want to briefly introduce the relevant concepts that are used within this work. First, the optical merit functions of interest will be specified. Subsequently, we will define the relevant measures from the AI world. With this framework parameterization we will investigate the dependency of the neural network performance for semantic segmentation on the optical quality of the perception chain. The non-linear correlation between those KPIs can be quantified by the Chatterjee’s rank correlation measure (Chatterjee,, 2021; Shi et al.,, 2021). We will shortly address the theoretical foundations of the Chatterjee’s rank correlation measure as it will be required in order to select the most suitable optical metric.
3.1 Optical merit functions
@@
Within this work, we are studying three different optical metrics, which have been used in previous work by Wolf et al., 2023c for the sensitivity analysis of AI-based algorithms for autonomous driving on optical wavefront aberrations induced by the windshield. These optical metrics can be evaluated if the wavefront aberration map is known a priori (Wolf et al., 2023c, ; Wolf et al., 2023b, ). Generally, the wavefront aberration map is decomposed into the orthogonal Zernike polynomial basis (Zernike and Stratton,, 1934; Zernike,, 1934; Bhatia and Wolf,, 1954) parameterized by the corresponding Zernike coefficients :
@@
From the optical path difference distribution across the aperture surface, parameterized by the normalized radial coordinate and the azimuth angle , the Point Spread Function (PSF) can be calculated by applying Fourier optical principles (Goodman,, 1968). In a nutshell, the PSF is the impulse response function or Green’s function of an optical system, which entirely determines the behavior of a Linear and Time-Invariant (LTI) system (Khoo,, 2018).
@@
The Fourier transform of the real-valued, incoherent PSF is known as the Optical Transfer Function (OTF). The OTF is generally complex-valued if the PSF is non-symmetric w.r.t. the optical axis. If the PSF is viewed as a scaled probability density function of the light distribution in the observer plane, then the PSF can be entirely characterized by its statistical moments (Bakker et al.,, 2008; Wolf et al., 2023c, ) and the OTF serves as the corresponding characteristic function. Consequently, the k-th order derivative of the OTF at zero spatial frequency entirely determines the k-th moment of the light distribution, e.g., gray values centroid (k=1), intensity variance (k=2) etc.
@@
The automotive industry is currently trying to grasp the importance of the OTF as an optical quality indicator function. Unfortunately, current attempts to map the OTF to a real-valued, scalar metric lead to insufficient optical KPIs as the Modulation Transfer Function (MTF) at half-Nyquist frequency (Wolf et al., 2023b, ; Mueller and Braun,, 2023). Analytically, the MTF is defined as the real part of the OTF (Goodman,, 1968).
@@
As an extension to mapping the OTF to a single spatial frequency value of the MTF, the Strehl ratio (Goodman,, 1968) has been proposed as an alternative measure to incorporate information about the entire spectrum into the mapping process. The Strehl ratio is defined as the spectral integral of the MTF in relation to the diffraction limited MTF area (Goodman,, 1968).
@@
An attempt to distill even more information into the mapping process of the OTF was made by Wolf et al., 2023c . They proposed the Optical Informative Gain (OIG) as the normalized spectral integral of the squared MTF function. This minor adjustment of the definition of the Strehl ratio is theoretically beneficial because the Strehl ratio exclusively captures information about the PSF at the optical axis. Hence, the captured information about the PSF, which entirely characterizes the optical system, is very limited and higher-order statistical moments are not accounted for at all. The OIG alleviates this situation by exploiting the Plancherel theorem (Deitmar and Echterhoff,, 2008) to retrieve information about the energy, which can be spatially discriminated in relation to the diffraction-limited case (Wolf et al., 2023c, ).
3.2 Neural network KPIs
@@
We will study the calibration quality of the pixel-wise confidences predicted by a CNN-based decoder head for semantic segmentation. The performance of the multi-class semantic segmentation task is evaluated by the mean Intersection over Union (mIoU) (Minaee et al.,, 2020):
@@
For assessing the calibration performance we will employ the mean ECE (mECE) as a point-wise calibration measure over all classes. The mECE, introduced by Pakdaman Naeini et al., (2015), is the weighted and binned average of the absolute difference between the model accuracy () and the confidence (). We will employ the softmax likelihood as a confidence estimator in alignment with the variation ratio (Maag et al.,, 2020). Mathematically, the mECE is given by:
@@
Finally, we define the Area Under the Reliability Error Curve (AUREC) as an additional point-wise calibration measure for classification, inspired by the UCS calibration metric for regression tasks (Wursthorn et al.,, 2024). The AUREC is equivalent to the mECE except that the weighting factor is set to one in order to artificially magnify the impact of predictions that fall into low confidence bins, which typically show a low cardinality (Wolf et al., 2024a, ).
3.3 Non-linear correlation measure
In order to quantify the correlation between two input signals, suppose and , the Pearson correlation coefficient is typically employed. Unfortunately, the Pearson correlation coefficient is restricted to linear relationships. In order to evaluate the non-linear correlation between the two signals – in our case ’optical quality’ and ’AI performance’ – alternative metrics are required. A very fundamental definition for non-linear correlation measures is given by the Dette-Siburg-Stoimenov’s rank correlation metric (Shi et al.,, 2021) defined as:
The quotient of the second term is given by the expected unexplained variance over the expected total variance. According to the law of variance decomposition (Weiss et al.,, 2005), the total variance is given as the sum of the explained variance and the unexplained variance . If there is a functional relationship between and then the unexplained variance vanishes. The expectation value of the unexplained variance of the indicator function over the distribution of is required in order to scan through all possible ranking thresholds according to their likelihood. Hence, the Dette-Siburg-Stoimenov’s correlation coefficient is a rank correlation metric.
Equation (4) is hard to evaluate numerically given a discrete sample. The Chatterjee’s rank correlation coefficient (Chatterjee,, 2021) presents an approximation of that converges to the expectation value as the sample size . The Chatterjee’s rank correlation coefficient is given by:
@@ 3.3 Non-linear correlation measure
4 PIPTS calibration architecture for semantic segmentation
@@
This work aims to demonstrate the benefits of incorporating physical priors to the PTS approach for confidence calibration. Since we are concerned with dataset shifts induced by optical aberrations of the windshield, our physical prior consists of the Zernike coefficient vector. Considering the dominant optical aberrations of windshields, we restrict the Zernike coefficient vector to the coefficients of the second radial order. Our AI target application in this study is semantic segmentation since a pixel-wise classification is supposed to be the most sensitive one w.r.t. the blurring operator. Generally, optical aberrations can be decomposed into two categories (Wolf et al., 2023b, ; Chan,, 2022): distortion (physically parameterized by the Zernike coefficients of the first radial order and described by the tilt operator) and blurring (mathematically expressed by the blurring operator and parameterized by the Zernike coefficients of radial order greater than one).
4.1 Baseline multi-task network
@@
In order to solve the target task and to provide the physical prior for the PIPTS network, we propose a CNN-based architecture with two coupled decoder heads and an additional residual encoder for identifying the Zernike coefficients of the second radial order (oblique astigmatism , defocus , orthogonal astigmatism ). The layout of the multi-task network is based on the UNET (Ronneberger et al.,, 2015) architecture and the predicted Zernike coefficients shall estimate the effective wavefront aberration map of the overall optical system consisting of windscreen and ADAS camera lens.
The network will be trained on the A2D2 (Geyer et al.,, 2020) dataset because it provides pixel-wise labels for high resolution images with pixels that were captured without a windshield. This is an essential requirement for the degradation model (see below), which is applied to enrich the data heterogeneity. Since we suspect telephoto cameras to be the most sensitive ones regarding optical aberrations induced by the windshield, we will only utilize the narrow-view, front-center camera (Sekonix SF3325-100 (Geyer et al.,, 2020)) of the A2D2 segmentation dataset. As a consequence, the training dataset comprises images and the test dataset consists of additional images, which corresponds to a test ratio of . Furthermore, the labeling taxonomy of the Cityscapes dataset (Cordts et al.,, 2016) was utilized to ensure the comparability of the presented results. For that reason, the ground truth annotations for the A2D2 classes were mapped to the classes Cityscapes taxonomy. The training of the neural network was performed on two Nvidia RTX A6000 with 48GB. The distributed training was terminated after the validation loss reached its minimum. To accomplish this, a learning rate schedule was applied that reduced the learning rate by a factor of if the validation loss did not improve during the last epochs and the training was finally terminated if the validation loss did not improve at all within the last epochs.
In order to enrich the A2D2 dataset with the optical aberrations induced by different windshield configurations, a Fourier optical degradation model (Wolf et al., 2023c, ; Goodman,, 1968) is employed for data augmentation. The range of the Zernike coefficients must be sufficiently sampled, such that the statistical complexity of the perception chain - individual part tolerances and installation tolerances - is accounted for. In this work, we will apply a uniform grid sampling of the Zernike coefficients of the second radial order in the range of . This does not guarantee that the complexity of the production process is sufficiently reflected, but it serves as a baseline for a proof of concept study, which is what this paper attempts to do.
@@
The detailed architecture of the baseline multi-task network is illustrated in Figure 3. The shared encoder consists of five encoder blocks with the number of filters doubling in each subsequent block. Each encoder block consists of two convolution layers followed by a batch normalization layer and a max pooling layer for downsampling. After the batch normalization layer, the computational graph is split into two branches to bypass information to the decoder for alleviating the vanishing gradient problem (He et al.,, 2016; Veit et al.,, 2016; Zaeemzadeh et al.,, 2018).
@@
The bottleneck of the UNET, distinguished by the lowest spatial feature resolution, consists of two convolution layers and subsequent batch normalization layers. After the bottleneck, the latent space representation is supposed to provide an embedding of the input information that optimally reflects the degrees of freedom of the underlying problem. Subsequently, the embedding is fed into two decoder heads. The decoder head for restoration aims to equalize the optical aberrations induced by the windshield and the decoder head for semantic segmentation targets on the pixel-wise classification.
@@
The decoder head for restoration consists of five transposed convolution blocks with the number of filters halving in each subsequent block. Each transposed convolution block consists of a transposed convolution layer followed by a batch normalization layer and a merging node in order to incorporate the high-fidelity information provided by the skip connection of the corresponding encoder block. The concatenated tensor is then smoothed by two subsequent convolution layers with stride equal to one, to preserve the dimensions.
@@
The decoder head for semantic segmentation is similar in structure to the decoder head for restoration, but the high-fidelity information from the encoder is merged with the corresponding restoration layer before entering the concatenation layer in the decoder. This decoder coupling is supposed to enhance the segmentation performance against optical aberrations.
@@
The downstream residual encoder for the Zernike coefficients identification consists of five ResNet (He et al.,, 2016) cells to classify the aberrations in terms of the Zernike coefficients of second order. Each ResNet cell consists of two convolution layers and two batch normalization layers, as well as a concatenation layer to merge the convolved signal with the input signal to alleviate the vanishing gradient problem (He et al.,, 2016; Veit et al.,, 2016; Zaeemzadeh et al.,, 2018). Finally, a max pooling layer is employed for downsampling. After the ResNet cells, the signal is flattened and fed into five dense layers with batch normalization and dropout for regularization.
@@
The multi-task network is trained by utilizing a customized loss function. The loss function consists of three components. The first component is the negative log-likelihood for semantic segmentation, aka. cross-entropy. The log-likelihood term is focused by an additional factor of , which enforces a focus of the learning process on classes that are hard to learn (Lin et al.,, 2017). Furthermore, class balancing is applied to equalize the representation of different classes within the dataset (Dreissig et al.,, 2023). The weighting factors for each class are calculated based on their occurrences in the dataset:
where denotes the number of instances for class and characterizes the total number of pixels within the A2D2 dataset. The second term quantifies the mismatch between the unperturbed image and the restored image by utilizing the L1-norm (Zhao et al.,, 2017). The third term quantifies the discrepancy between the predicted Zernike coefficient vector and the ground truth vector by applying the L2-norm. All components are summed up by considering individual weighting factors, which are determined during hypertuning.
@@
In addition to the weighting factor for the restoration loss and the Zernike loss, there are four other hyperparameters to tune:
@@
The weighting factor of the Kernel-Orthonormality-Regularizer (KOR) term (see below), which is added linearly to the loss function as a penalty.
@@
The batchsize, which relates to the number of images processed before the trainable variables are updated.
When employing a large batchsize, the model’s quality often degrades, particularly in terms of its generalization capabilities. Models with large batchsizes are prone to reaching sharp minima in the loss landscape, which are generally associated with reduced generalization performance. Conversely, small batchsizes tend to converge to flatter minima due to the inherent noise in the gradient estimation (Keskar et al.,, 2017).
For the multi-task network, as for the PIPTS model, the input and output is normalized to zero mean and unit variance in order to equalize the dissimilarity of feature units.
Generally, the Gaussian Error Linear Unit (GELU) (Hendrycks and Gimpel,, 2023) is used as an activation function for the majority of layers. GELU is defined as , where is the standard Gaussian cumulative distribution function. The GELU activation function returns a likelihood-based output, unlike the Rectified Linear Unit (ReLU) (Nair and Hinton,, 2010), which simply gates the input according to the sign. Generalizing from monotonic (e.g. ReLU) to non-monotonic (e.g. GELU) activations can increase a neuron’s discriminative capacity as has been demonstrated for the XOR problem (Bauckhage and Speicher,, 2019).
In order to enhance the generalization capabilities of the model, Kernel Orthonormality Regularization (KOR) (Kim and Yun,, 2022) is utilized. KOR penalizes orthogonality violations of the convolutional kernel matrices leading to reduced feature redundancy, which enriches the information capacity of the latent space embedding and boosts the model generalizability. To implement this, the convolutional kernel tensor is reshaped to a 2D-kernel matrix maintaining the innermost dimension (number of output channels). Afterwards, the Gramian matrix is computed from the kernel matrix and the Frobenius norm is used for quantifying the residuals w.r.t. the identity matrix. The Frobenius norm corresponds to the Euclidean norm of the vector of eigenvalues of the matrix.
4.2 PIPTS calibration network

### v2

@@ III Theoretical essentials
In this section, we want to briefly introduce the relevant concepts that are used within this work. First, the optical merit functions of interest will be specified. Subsequently, we will define the relevant measures from the AI world. With this framework parameterization we will investigate the dependency of the neural network performance for semantic segmentation on the optical quality of the perception chain. The non-linear correlation between those KPIs can be quantified by the Chatterjee’s rank correlation measure [35, 36]. We will shortly address the theoretical foundations of the Chatterjee’s rank correlation measure as it will be required in order to select the most suitable optical metric.
III-A Optical merit functions
@@
Within this work, we are studying three different optical metrics, which have been used in previous work by Wolf et al. [5] for the sensitivity analysis of AI-based algorithms for autonomous driving on optical wavefront aberrations induced by the windshield. These optical metrics can be evaluated if the wavefront aberration map is known a priori [5, 37]. Generally, the wavefront aberration map is decomposed into the orthogonal Zernike polynomial basis [26, 27, 28] parameterized by the corresponding Zernike coefficients :
@@
From the optical path difference distribution across the aperture surface, parameterized by the normalized radial coordinate and the azimuth angle , the Point Spread Function (PSF) can be calculated by applying Fourier optical principles [38]. In a nutshell, the PSF is the impulse response function or Green’s function of an optical system, which entirely determines the behavior of a Linear and Time-Invariant (LTI) system [39].
@@
The Fourier transform of the real-valued, incoherent PSF is known as the Optical Transfer Function (OTF). The OTF is generally complex-valued if the PSF is non-symmetric w.r.t. the optical axis. If the PSF is viewed as a scaled probability density function of the light distribution in the observer plane, then the PSF can be entirely characterized by its statistical moments [40, 5] and the OTF serves as the corresponding characteristic function. Consequently, the k-th order derivative of the OTF at zero spatial frequency entirely determines the k-th moment of the light distribution, e.g., gray values centroid (k=1), intensity variance (k=2) etc.
@@
The automotive industry is currently trying to grasp the importance of the OTF as an optical quality indicator function. Unfortunately, current attempts to map the OTF to a real-valued, scalar metric lead to insufficient optical KPIs as the Modulation Transfer Function (MTF) at half-Nyquist frequency [37, 41]. Analytically, the MTF is defined as the real part of the OTF [38].
@@
As an extension to mapping the OTF to a single spatial frequency value of the MTF, the Strehl ratio [38] has been proposed as an alternative measure to incorporate information about the entire spectrum into the mapping process. The Strehl ratio is defined as the spectral integral of the MTF in relation to the diffraction limited MTF area [38].
@@
An attempt to distill even more information into the mapping process of the OTF was made by Wolf et al. [5]. They proposed the Optical Informative Gain (OIG) as the normalized spectral integral of the squared MTF function. This minor adjustment of the definition of the Strehl ratio is theoretically beneficial because the Strehl ratio exclusively captures information about the PSF at the optical axis. Hence, the captured information about the PSF, which entirely characterizes the optical system, is very limited and higher-order statistical moments are not accounted for at all. The OIG alleviates this situation by exploiting the Plancherel theorem [42] to retrieve information about the energy, which can be spatially discriminated in relation to the diffraction-limited case [5].
III-B Neural network KPIs
@@
We will study the calibration quality of the pixel-wise confidences predicted by a CNN-based decoder head for semantic segmentation. The performance of the multi-class semantic segmentation task is evaluated by the mean Intersection over Union (mIoU) [43]:
@@
For assessing the calibration performance we will employ the mean ECE (mECE) as a point-wise calibration measure over all classes. The mECE, introduced by Naeini et al. [6], is the weighted and binned average of the absolute difference between the model accuracy () and the confidence (). We will employ the softmax likelihood as a confidence estimator in alignment with the variation ratio [9]. Mathematically, the mECE is given by:
@@
Finally, we define the Area Under the Reliability Error Curve (AUREC) as an additional point-wise calibration measure for classification, inspired by the UCS calibration metric for regression tasks [44]. The AUREC is equivalent to the mECE except that the weighting factor is set to one in order to artificially magnify the impact of predictions that fall into low confidence bins, which typically show a low cardinality [10].
III-C Non-linear correlation measure
In order to quantify the correlation between two input signals, suppose and , the Pearson correlation coefficient is typically employed. Unfortunately, the Pearson correlation coefficient is restricted to linear relationships. In order to evaluate the non-linear correlation between the two signals – in our case ’optical quality’ and ’AI performance’ – alternative metrics are required. A very fundamental definition for non-linear correlation measures is given by the Dette-Siburg-Stoimenov’s rank correlation metric [36] defined as:
The quotient of the second term is given by the expected unexplained variance over the expected total variance. According to the law of variance decomposition [45], the total variance is given as the sum of the explained variance and the unexplained variance . If there is a functional relationship between and then the unexplained variance vanishes. The expectation value of the unexplained variance of the indicator function over the distribution of is required in order to scan through all possible ranking thresholds according to their likelihood. Hence, the Dette-Siburg-Stoimenov’s correlation coefficient is a rank correlation metric.
Equation (4) is hard to evaluate numerically given a discrete sample. The Chatterjee’s rank correlation coefficient [35] presents an approximation of that converges to the expectation value as the sample size . The Chatterjee’s rank correlation coefficient is given by:
@@ III-C Non-linear correlation measure
IV PIPTS calibration architecture for semantic segmentation
@@
This work aims to demonstrate the benefits of incorporating physical priors into the PTS approach for confidence calibration. Since we are concerned with dataset shifts induced by optical aberrations of the windshield, our physical prior consists of the Zernike coefficient vector. In this paper, semantic segmentation is selected as the AI target application due to its intrinsic sensitivity to the impact of the blurring operator, as it is based on pixel-level classifications.
IV-A Physical priors
Generally, optical aberrations can be decomposed into distortion and blurring [37, 46]. The tilt operator describes the impact of distortions on the imaging process and is parameterized by the Zernike coefficients of the first radial order. On the other hand, the blurring operator is parameterized by the Zernike coefficients of radial order greater than one. In this study, we will solely focus on the impact of the blurring operator on the imaging process, as a measurement procedure to correct for distortions is already available [47]. Considering the dominant optical aberrations of windshields, we restrict the Zernike coefficient vector to the coefficients of the second radial order, which includes oblique astigmatism (), defocus (), and orthogonal astigmatism (). To model the effect of these aberrations, we apply a Fourier optical degradation model [5, 38] for data augmentation. The Zernike coefficients are sampled uniformly in the range . This does not guarantee that the complexity of the production process is sufficiently reflected, but it serves as a physically sound baseline for our proof of concept study. For a comprehensive modeling of the statistical complexity of the perception chain’s manufacturing process, individual part tolerances and installation tolerances need to be accounted for.
IV-B Architecture of the baseline multi-task network
@@
In order to solve the target task and to provide the physical prior for the PIPTS calibration network, we propose a CNN-based multi-task architecture with two coupled decoder heads and an additional residual encoder for estimating the Zernike coefficients of the second radial order of the effective wavefront aberration map of the overall optical system consisting of windscreen and ADAS camera lens. The detailed architecture of the baseline multi-task network is illustrated in Figure 3.
@@
The shared encoder consists of five encoder blocks, each with two convolution layers, two batch normalization layers, and a max pooling layer for downsampling. The number of filters is doubled in each successive encoder block and after the last batch normalization layer, the computational graph is split into two branches to bypass information to the decoder for alleviating the vanishing gradient problem [48, 49, 50].
@@
The bottleneck of the UNET, distinguished by the lowest spatial feature resolution, consists of two convolution layers and subsequent batch normalization layers. After the bottleneck, the latent space representation is supposed to provide an embedding of the input information that optimally reflects the degrees of freedom of the underlying problem. Subsequently, the embedding is fed into two decoder heads.
@@
The restoration decoder aims to reverse the degradation induced by the windshield and mirrors the encoder with five transposed convolution blocks, each with a transposed convolution layer, three batch normalization layers, and a merging node that incorporates high-fidelity information from the encoder. The number of filters is halved in each successive decoder block and the concatenated tensor is smoothed by two convolution layers with stride equal to one.
@@
The segmentation decoder targets on the pixel-wise classification and also consists of five transposed convolution blocks. The high-fidelity information from the encoder is merged with the corresponding restoration layer before entering the concatenation layer in the decoder. This decoder coupling is supposed to enhance the segmentation performance against optical aberrations.
@@
The downstream residual encoder for the regression of the Zernike coefficients comprises five ResNet [48] cells. Each cell contains two convolution layers and two batch normalization layers, as well as a residual connection to merge the convolved signal with the input signal in order to alleviate the vanishing gradient problem [48, 49, 50]. Finally, a max pooling layer is employed for downsampling. Subsequent to the ResNet cells, the signal is flattened and passed through five dense layers with batch normalization and dropout for regularization.
IV-C Dataset and preprocessing
The network will be trained on the A2D2 [51] dataset because it provides pixel-wise labels for high resolution images with pixels that were captured without a windshield, a necessary condition for generating synthetically degraded images in order to enrich the data heterogeneity. Only the images from the narrow-view front-center camera (Sekonix SF3325-100 [51]) are used due to its sensitivity to optical aberrations. The dataset comprises training images and test images, corresponding to a test split. To ensure comparability, the -class ground truth annotations of the A2D2 [51] dataset are mapped to the -class labeling taxonomy of the Cityscapes [52] dataset. For the multi-task network, as for the PIPTS calibration network, the input and output tensors are normalized to zero mean and unit variance in order to equalize the dissimilarity of feature units.
IV-D Training configuration
The training was distributed across two Nvidia RTX A6000 GPUs with 48GB memory. A learning rate schedule reduced the learning rate by a factor of if the validation loss plateaued for the last epochs, and the training was terminated if no improvement occurred over the last epochs.
IV-E Loss Function
The multi-task network is trained by utilizing a customized loss function with three components:
1.
@@ IV-E Loss Function
Semantic segmentation loss: Negative log-likelihood (aka. cross-entropy) loss with a focal modulation term , which enforces a focus of the learning process on classes that are hard to learn [53]. Additionally, class balancing is applied to equalize the representation of different classes within the dataset [7]. The weighting factors for each class are calculated based on their frequencies in the dataset:
(7) where denotes the number of instances for class , characterizes the total number of pixels within the A2D2 dataset, and specifies the number of classes.
2.
Restoration loss: L1-norm between the unperturbed images and the restored images [54].
3.
Zernike regression loss: L2-norm between the predicted and the ground truth Zernike coefficient vectors.
The restoration loss and the Zernike regression loss add to the total loss by considering an associated weighting factor, determined via hyperparameter tuning.
IV-F Hyperparameters
@@
In addition to the loss term weighting factors, there are four other hyperparameters to tune:
1.
@@
The weighting factor of the Kernel-Orthonormality-Regularizer (KOR) term (see Section IV-H), which is added linearly to the loss function as a penalty.
2.
@@
4.
The batchsize, which relates to the number of images processed before the trainable variables are updated. When employing a large batchsize, the model’s quality often degrades, particularly in terms of its generalization capabilities. Models with large batchsizes are prone to reaching sharp minima in the loss landscape, which are generally associated with reduced generalization performance. Conversely, small batchsizes tend to converge to flatter minima due to the stochasticity induced into the gradient estimates [55].
IV-G Activation functions
The Gaussian Error Linear Unit (GELU) [56] is used as an activation function for the shared encoder, the segmentation decoder, and the restoration decoder. GELU is defined as , where is the standard Gaussian cumulative distribution function. The GELU activation function returns a likelihood-based output, unlike the Rectified Linear Unit (ReLU) [57], which simply gates the input according to the sign. Generalizing from monotonic (e.g. ReLU) to non-monotonic (e.g. GELU) activation functions can increase a neuron’s discriminative capacity as has been demonstrated for the XOR problem [58]. For the regression task, the Exponential Linear Unit (ELU) [59] is employed in the ResNet cells and the dense layers. Moreover, the activation functions used in the output layers are specifically tailored for their respective tasks. The final layer of the segmentation head applies a softmax activation function. The restoration decoder implements the ReLU as an activation function for the output layer. Furthermore, the regression output is obtained by a linear activation function.
IV-H Model regularization
In order to enhance the generalization capabilities of the model, Kernel-Orthonormality-Regularization (KOR) [60] is utilized. KOR penalizes orthonormality violations of the convolutional kernel matrices leading to reduced feature redundancy, which enriches the information capacity of the latent space embedding and boosts the model generalizability. To implement this, the convolutional kernel tensor is reshaped to a 2D-kernel matrix maintaining the innermost dimension (number of output channels). Afterwards, the Gramian matrix is computed from the kernel matrix and the Frobenius norm is used for quantifying the residuals w.r.t. the identity matrix. The Frobenius norm corresponds to the Euclidean norm of the vector of eigenvalues of the matrix.
IV-I PIPTS calibration network